OpenStudy (anonymous):

I am having trouble factoring this (a^3x-b^x)-(a^3y+b^3y) Can anyone help ?

OpenStudy (bahrom7893):

difference of cubes - sum of cubes, factor each separately

OpenStudy (anonymous):

this IS NOT difference of cubes

OpenStudy (anonymous):

I started by Factoring out x and y X(a^3-b^3)-y(a^3+b^3)

OpenStudy (anonymous):

@precal

OpenStudy (eyust707):

\[a^3 + b^3 = (a + b)(a^2 – ab + b^2) \]

OpenStudy (eyust707):

\[a^3 – b^3 = (a – b)(a^2 + ab + b^2)\]

OpenStudy (anonymous):

thats not what it says

OpenStudy (eyust707):

typo?

OpenStudy (anonymous):

So X(a-b)(a^2+ab+b^2)-y(a+b)(a^2-ab+b^2) ?

OpenStudy (anonymous):

if it is this \[(a^3 x-b^3 x)-(a^3 y+b^3 y)\] you cannot factor it

OpenStudy (anonymous):

you could rewrite it as \[x(a-b)(a^2+ab+b^2)-y(a+b)(a^2-ab+b^2)\] but that is not factoring

OpenStudy (anonymous):

factor means to write as a product, and this is still a difference (minus sign in the middle)

OpenStudy (precal):

 you need to write problem as given, the parenthesis changes the problem altogether